given f(x)=5x^2-3x+1,find f(-2)
I do not understand this-functions are not clear to me

Answer by rajagopalan(174)

given f(x)=5x^2-3x+1,find f(-2)
I do not understand this-functions are not clear to me
Answer:
f(x)=5x^2-3x+1,
f(-2)= 5(-2^2)-(3*-2)+1 just substitute -2 in place of x
=5*(-2*-2)-(-6)+1
=5*4+6+1
=20+6-1
=26-1
=25.
Okay !